# Integrative Learning of Intensity Fluctuations of Quantum Dots under Excitation via a Tailored Mixture Hidden Markov Model

### v1

@@ 4.4 Simulation Results
The simulation results, presented in Tables 2 and 3, evaluate the performance of the proposed MHMM-$\beta$ model and its close competitor under balanced and unbalanced cluster settings, across varying sample sizes ($N$) and time series lengths ($T$). Also, the oracle MHMM-$\beta^{*}$ bypasses the clustering step and achieves the lowest parameter estimation errors, which serves as a benchmark for the achievable accuracy.
In the balanced cluster setting (Table 2), the proposed MHMM-$\beta$ consistently outperforms the competing method in terms of clustering accuracy (CC), achieving rates close to or exceeding 96% across all scenarios. The clustering performance of MHMM-G is significantly lower, with CC values around 30-36%, underscoring the benefit of using the inflated Beta distribution for modeling bounded intensities with inflation at 0 and 1. The estimation errors of MHMM-$\beta$ are only slightly higher compared to the oracle method MHMM-$\beta^{*}$, as expected. Increasing the sample size ($N$) and time series length ($T$) generally leads to improved parameter estimation, reflecting the efficiency of the proposed method in utilizing larger data.
In the unbalanced cluster setting (Table 3), the clustering accuracy (CC) of MHMM-$\beta$ remains strong, ranging from 81% to 85%. The estimation errors for both transition matrices and state parameters are slightly higher than in the balanced setting. Nonetheless, the MHMM-$\beta$ model consistently outperforms MHMM-G.
These results underscore the effectiveness of the MHMM-$\beta$ model in accurately capturing both the state dynamics and the clustering structure of the QD intensities, making it a powerful tool for integrative analysis in complex scenarios.
Table 2: Simulation: Results under the balanced settings.
@@
Method Length ($T$) $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC
Sample size ($N$) = 50
MHMM-$\beta^{*}$ 1000 0.56 (0.03) 0.17 (0.00)
MHMM-G 0.98 (0.01) 0.36 (0.01)
MHMM-$\beta$ 1.56 (0.11) 0.32 (0.02) 0.96 (0.01)
Sample size ($N$) = 50
MHMM-$\beta^{*}$ 2000 0.52 (0.04) 0.17 (0.00)
MHMM-G 0.97 (0.01) 0.31 (0.00)
MHMM-$\beta$ 1.66 (0.13) 0.28 (0.02) 0.98 (0.01)
Sample size ($N$) = 100
MHMM-$\beta^{*}$ 1000 0.54 (0.03) 0.17 (0.00)
MHMM-G 0.97 (0.01) 0.33 (0.00)
MHMM-$\beta$ 1.46 (0.13) 0.26 (0.02) 0.98 (0.01)
Sample size ($N$) = 100
MHMM-$\beta^{*}$ 2000 0.50 (0.03) 0.16 (0.00)
MHMM-G 0.97 (0.01) 0.31 (0.00)
MHMM-$\beta$ 1.48 (0.12) 0.27 (0.02) 0.98 (0.01)
Table 3: Simulation: Results under the unbalanced settings.
@@
Method Length ($T$) $\mbox{Er}(\widehat{\boldsymbol{\Theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC
Sample size ($N$) = 50
MHMM-$\beta^{*}$ 1000 0.78 (0.04) 0.18 (0.00)
MHMM-G 1.22 (0.02) 0.59 (0.01)
MHMM-$\beta$ 1.08 (0.08) 0.59 (0.02) 0.81 (0.01)
Sample size ($N$) = 50
MHMM-$\beta^{*}$ 2000 0.67 (0.03) 0.17 (0.00)
MHMM-G 1.24 (0.02) 0.57 (0.01)
MHMM-$\beta$ 1.02 (0.08) 0.54 (0.03) 0.85 (0.01)
Sample size ($N$) = 100
MHMM-$\beta^{*}$ 1000 0.77 (0.04) 0.17 (0.00)
MHMM-G 1.21 (0.02) 0.59 (0.01)
MHMM-$\beta$ 1.25 (0.11) 0.58 (0.03) 0.82 (0.02)
Sample size ($N$) = 100
MHMM-$\beta^{*}$ 2000 0.74 (0.04) 0.16 (0.00)
MHMM-G 1.22 (0.02) 0.58 (0.01)
MHMM-$\beta$ 1.16 (0.12) 0.56 (0.03) 0.84 (0.02)
5 Case Study: Integrative Analysis of Multiple Quantum Dots
@@ 5.1 Cluster Patterns
We applied the proposed MHMM-$\beta$ approach with varying number of clusters. The cluster patterns are displayed in Figure 4, in which the colors or the grey levels represent different clusters and the stripes show how the dots are put into more and more clusters from the left to the right.
@@
Starting from the 2-cluster model, it can be seen that in the 3-cluster model the third cluster is formed by QDs from both clusters. From then on, when the number of clusters increases, the new clusters are mainly from further splitting the third cluster, and the first two clusters remain relatively stable throughout. This clear pattern suggests that the 3-cluster model is the most stable and informative. We have also computed BIC and ICL values for these models and unfortunately they had a big discrepancy: BIC would select the 6-cluster model while ICC suggested 1 or 2 clusters. This in fact is consistent with comparative studies in the literature, which have shown that when the clusters are too overlapped, ICL tends to favor less number of clusters than BIC. From these results and after consulting with our chemistry collaborators, we decided to mainly focus on the results from the three-cluster model; as to be shown below, the results indeed are highly interpretable based on principles of chemistry and physics. We also provide the results from other models in Section B of the Supplementary Material.
5.2 Results from 3-Cluster Model
In Figure 5, the right panel shows the estimated 0/1 inflated Beta distributions of the three intensity states from the MHMM-$\beta$ model. With the estimated state series, we also computed the empirical distributions of the three states directly from the standardized intensity data, which are shown in the left panel. As expected, the estimated and the empirical distributions are quite similar to each other, suggesting that here are no apparent outliers and the MHMM-$\beta$ model fits the data well and successfully captured the collective behaviors of intensity fluctuations of all QDs.

### v2

@@ 4.4 Simulation Results
Tables 25 present simulation results under two different scenarios, each tested with balanced and unbalanced cluster partitions. The oracle procedure, MHMM-$\beta^{*}$, provides a benchmark for parameter estimation by assuming that the cluster memberships are known and the true estimates are provided as initials. As expected, it achieves the lowest estimation errors, offering insight into how well the proposed method and its competitors perform.
Across all simulation settings, the proposed MHMM-$\beta$ method performs well in parameter estimation and clustering accuracy. In many cases, the estimation errors for MHMM-$\beta$ come close to those attained by the oracle method. In both balanced and unbalanced scenarios, MHMM-$\beta$ outperforms the single-dot methods (HMM-G and HMM-$\beta$), which do not pool information across QDs. These single-dot approaches tend to have higher errors in transition matrix estimation and yield lower accuracy in retrieving the true cluster structure. The MHMM-G method partially captures some shared dynamics among QDs but is also consistently outperformed by MHMM-$\beta$. This difference emphasizes the value of an inflated-Beta emission distribution for modeling the standardized intensities and handling zero/one inflation features that cannot be adequately addressed by Gaussian assumptions.
When comparing Scenario 1 vs. 2 and balanced vs. unbalanced partitions, the settings of more overlapping clusters and unbalanced partitions pose more challenges in identifying and estimating smaller clusters, leading to slightly higher errors and lower clustering accuracy; nonetheless, MHMM-$\beta$ remains robust and still outperforms its competitors under these conditions. Across all methods, longer time series contribute to more accurate parameter estimates and higher clustering accuracy.
Overall, the simulation studies confirm that MHMM-$\beta$ successfully captures both the hidden-state dynamics and the QD clustering patterns. By aligning the state/emission distributions with the physical and experimental characteristics of the QDs, the proposed method offers considerable advantages in parameter estimation and clustering.
Additional simulation results are reported in Section C of the Supplemental Materials. In particular, we have conducted simulation studies to examine the performance of several information criteria, including the Akaike Information Criterion (AIC) (Akaike, 1974), the Bayesian Information Criterion (BIC) (Schwarz, 1978), and the Integrated Completed Likelihood (ICL) criterion (Biernacki et al., 2000), on selecting the number of clusters. Our results illustrate the well-known differences in how each criterion penalizes model complexity and cluster separation; while they often converge on the truth, discrepancies sometimes arise in cases of overlapping or unevenly sized clusters. Therefore, in practice, we advocate the examination of the cluster patterns of different models and the use of both information criteria and domain knowledge for model selection.
Table 2: Simulation: Results for Scenario 1 with balanced partition. For better presentation, values in $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ and $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ are multiplied by $10$, values in $\mbox{Er}(\widehat{\mu})$ are multiplied by $10^{2}$, and values in $\mbox{Er}(\widehat{\sigma}^{2})$ are multiplied by $10^{3}$.
@@
Method Length ($T$) $\mbox{Er}(\widehat{{\mu}})$ $\mbox{Er}(\widehat{\sigma}^{2})$ $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC Balanced, Sample size ($N$) = 100 MHMM-$\beta^{*}$ 250 0.36 (0.02) 0.77 (0.05) 0.00 (0.00) 3.17 (0.13) 0.13 (0.01) HMM-G 14.37 (0.11) 9.95 (0.06) 3.21 (0.09) 2.07 (0.02) 0.53 (0.00) HMM-$\beta$ 20.61 (0.15) 9.73 (0.08) 3.78 (0.07) 2.15 (0.03) 0.53 (0.00) MHMM-G 0.74 (0.03) 2.19 (0.08) 1.13 (0.13) 0.62 (0.06) 0.89 (0.01) MHMM-$\beta$ 0.38 (0.02) 0.79 (0.05) 0.60 (0.11) 3.26 (0.13) 0.36 (0.05) 0.94 (0.01) MHMM-$\beta^{*}$ 500 0.27 (0.01) 0.53 (0.03) 0.00 (0.00) 2.08 (0.09) 0.11 (0.01) HMM-G 13.68 (0.09) 8.35 (0.04) 3.72 (0.05) 2.08 (0.02) 0.52 (0.00) HMM-$\beta$ 19.88 (0.13) 6.73 (0.06) 3.74 (0.05) 1.99 (0.02) 0.53 (0.00) MHMM-G 0.78 (0.03) 2.25 (0.06) 1.70 (0.16) 0.83 (0.07) 0.86 (0.01) MHMM-$\beta$ 0.28 (0.01) 0.55 (0.03) 0.80 (0.15) 2.36 (0.10) 0.42 (0.06) 0.93 (0.01) MHMM-$\beta^{*}$ 1000 0.19 (0.01) 0.40 (0.02) 0.00 (0.00) 1.58 (0.06) 0.11 (0.01) HMM-G 12.79 (0.08) 7.51 (0.04) 3.60 (0.03) 2.03 (0.01) 0.57 (0.00) HMM-$\beta$ 18.69 (0.11) 5.11 (0.05) 3.79 (0.02) 1.95 (0.01) 0.58 (0.00) MHMM-G 0.65 (0.02) 2.16 (0.04) 1.56 (0.17) 0.70 (0.06) 0.87 (0.01) MHMM-$\beta$ 0.22 (0.01) 0.43 (0.02) 0.58 (0.14) 1.84 (0.08) 0.34 (0.05) 0.96 (0.01) MHMM-$\beta^{*}$ 2000 0.14 (0.01) 0.26 (0.01) 0.00 (0.00) 1.12 (0.04) 0.09 (0.01) HMM-G 11.81 (0.04) 7.18 (0.02) 3.71 (0.04) 2.01 (0.02) 0.61 (0.01) HMM-$\beta$ 17.73 (0.07) 4.25 (0.03) 3.91 (0.01) 1.95 (0.01) 0.61 (0.00) MHMM-G 0.67 (0.02) 2.18 (0.03) 1.25 (0.16) 0.61 (0.06) 0.91 (0.01) MHMM-$\beta$ 0.16 (0.01) 0.35 (0.02) 0.69 (0.15) 1.51 (0.07) 0.43 (0.06) 0.95 (0.01)
Table 3: Simulation: Results for Scenario 1 with unbalanced partition. For better presentation, values in $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ and $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ are multiplied by $10$, values in $\mbox{Er}(\widehat{\mu})$ are multiplied by $10^{2}$, and values in $\mbox{Er}(\widehat{\sigma}^{2})$ are multiplied by $10^{3}$.
@@
Method Length ($T$) $\mbox{Er}(\widehat{{\mu}})$ $\mbox{Er}(\widehat{\sigma}^{2})$ $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC Unbalanced, Sample size ($N$) = 100 MHMM-$\beta^{*}$ 250 0.37 (0.02) 0.66 (0.04) 0.00 (0.00) 3.35 (0.15) 0.13 (0.01) HMM-G 8.33 (0.09) 8.26 (0.06) 1.34 (0.08) 2.08 (0.02) 0.76 (0.01) HMM-$\beta$ 14.62 (0.14) 7.12 (0.08) 1.25 (0.04) 2.20 (0.02) 0.80 (0.00) MHMM-G 1.25 (0.03) 3.46 (0.06) 1.74 (0.13) 1.10 (0.07) 0.84 (0.01) MHMM-$\beta$ 0.38 (0.02) 0.69 (0.04) 1.92 (0.16) 3.74 (0.15) 0.49 (0.05) 0.85 (0.01) MHMM-$\beta^{*}$ 500 0.26 (0.01) 0.48 (0.03) 0.00 (0.00) 2.24 (0.10) 0.12 (0.01) HMM-G 7.97 (0.05) 7.42 (0.04) 1.31 (0.02) 2.15 (0.02) 0.82 (0.00) HMM-$\beta$ 13.98 (0.12) 4.71 (0.06) 1.45 (0.03) 2.17 (0.02) 0.81 (0.00) MHMM-G 1.14 (0.03) 3.39 (0.06) 1.15 (0.11) 0.93 (0.07) 0.89 (0.01) MHMM-$\beta$ 0.27 (0.01) 0.53 (0.03) 0.98 (0.14) 2.63 (0.13) 0.38 (0.05) 0.93 (0.01) MHMM-$\beta^{*}$ 1000 0.21 (0.01) 0.39 (0.02) 0.00 (0.00) 1.69 (0.06) 0.14 (0.02) HMM-G 7.76 (0.04) 7.01 (0.03) 1.30 (0.02) 2.19 (0.02) 0.83 (0.00) HMM-$\beta$ 13.43 (0.09) 3.57 (0.04) 1.39 (0.02) 2.17 (0.02) 0.83 (0.00) MHMM-G 1.11 (0.02) 3.38 (0.05) 1.04 (0.10) 0.77 (0.06) 0.91 (0.01) MHMM-$\beta$ 0.22 (0.01) 0.43 (0.02) 0.47 (0.10) 2.15 (0.11) 0.27 (0.03) 0.97 (0.01) MHMM-$\beta^{*}$ 2000 0.13 (0.01) 0.26 (0.01) 0.00 (0.00) 1.14 (0.05) 0.12 (0.01) HMM-G 7.63 (0.03) 6.79 (0.02) 1.22 (0.02) 2.17 (0.02) 0.84 (0.00) HMM-$\beta$ 12.97 (0.07) 3.08 (0.03) 1.33 (0.01) 2.17 (0.01) 0.84 (0.00) MHMM-G 1.07 (0.02) 3.37 (0.03) 0.64 (0.11) 0.45 (0.04) 0.95 (0.01) MHMM-$\beta$ 0.15 (0.01) 0.39 (0.02) 0.24 (0.08) 2.21 (0.10) 0.20 (0.02) 0.98 (0.01)
Table 4: Simulation: Results for Scenario 2 with balanced partition. For better presentation, values in $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ and $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ are multiplied by $10$, and values in $\mbox{Er}(\widehat{\mu})$ and $\mbox{Er}(\widehat{\sigma}^{2})$ are multiplied by $10^{2}$.
Method Length ($T$) $\mbox{Er}(\widehat{{\mu}})$ $\mbox{Er}(\widehat{\sigma}^{2})$ $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC Balanced, Sample size ($N$) = 100 MHMM-$\beta^{*}$ 250 0.58 (0.03) 2.25 (0.11) 0.00 (0.00) 3.31 (0.13) 0.10 (0.00) HMM-G 6.52 (0.04) 32.94 (0.05) 1.05 (0.06) 0.98 (0.01) 0.75 (0.01) HMM-$\beta$ 5.81 (0.26) 9.01 (0.18) 1.43 (0.09) 0.78 (0.04) 0.73 (0.01) MHMM-G 6.79 (0.08) 30.73 (0.09) 0.79 (0.07) 1.00 (0.02) 0.79 (0.01) MHMM-$\beta$ 0.92 (0.05) 2.92 (0.15) 0.38 (0.04) 7.95 (0.51) 0.22 (0.01) 0.97 (0.00) MHMM-$\beta^{*}$ 500 0.40 (0.02) 1.54 (0.06) 0.00 (0.00) 2.27 (0.09) 0.07 (0.00) HMM-G 6.65 (0.03) 31.94 (0.03) 0.86 (0.06) 0.89 (0.01) 0.83 (0.00) HMM-$\beta$ 2.86 (0.09) 6.84 (0.16) 0.50 (0.03) 0.32 (0.01) 0.90 (0.00) MHMM-G 6.71 (0.07) 30.75 (0.07) 0.57 (0.07) 0.94 (0.01) 0.88 (0.01) MHMM-$\beta$ 0.70 (0.04) 2.21 (0.12) 0.21 (0.08) 6.28 (0.39) 0.19 (0.02) 0.98 (0.01) MHMM-$\beta^{*}$ 1000 0.28 (0.01) 1.03 (0.05) 0.00 (0.00) 1.73 (0.08) 0.05 (0.00) HMM-G 6.72 (0.02) 31.35 (0.03) 0.83 (0.06) 0.88 (0.01) 0.90 (0.00) HMM-$\beta$ 1.60 (0.05) 4.95 (0.11) 0.35 (0.02) 0.23 (0.00) 0.96 (0.00) MHMM-G 6.54 (0.06) 30.79 (0.07) 0.25 (0.05) 0.91 (0.01) 0.97 (0.01) MHMM-$\beta$ 0.84 (0.07) 2.31 (0.17) 0.01 (0.01) 7.70 (0.64) 0.16 (0.01) 1.00 (0.00) MHMM-$\beta^{*}$ 2000 0.19 (0.01) 0.75 (0.03) 0.00 (0.00) 1.21 (0.05) 0.04 (0.00) HMM-G 6.79 (0.01) 30.97 (0.02) 0.54 (0.03) 0.89 (0.00) 0.96 (0.00) HMM-$\beta$ 1.00 (0.02) 3.96 (0.09) 0.07 (0.01) 0.21 (0.00) 0.99 (0.00) MHMM-G 6.50 (0.06) 30.68 (0.07) 0.02 (0.01) 0.88 (0.01) 1.00 (0.00) MHMM-$\beta$ 0.59 (0.04) 1.78 (0.13) 0.08 (0.06) 5.70 (0.48) 0.15 (0.02) 0.99 (0.01)
Table 5: Simulation: Results for Scenario 2 with unbalanced partition. For better presentation, values in $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ and $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ are multiplied by $10$, and values in $\mbox{Er}(\widehat{\mu})$ and $\mbox{Er}(\widehat{\sigma}^{2})$ are multiplied by $10^{2}$.
Method Length ($T$) $\mbox{Er}(\widehat{{\mu}})$ $\mbox{Er}(\widehat{\sigma}^{2})$ $\mbox{Er}(\widehat{\boldsymbol{\delta}})$ $\mbox{Er}(\widehat{\boldsymbol{\theta}})$ $\mbox{Er}(\widehat{\boldsymbol{\Pi}})$ CC Unbalanced, Sample size ($N$) = 100 MHMM-$\beta^{*}$ 250 0.69 (0.03) 2.64 (0.11) 0.00 (0.00) 3.51 (0.17) 0.12 (0.00) HMM-G 6.07 (0.03) 33.44 (0.04) 2.05 (0.13) 1.24 (0.03) 0.74 (0.01) HMM-$\beta$ 4.88 (0.24) 11.06 (0.25) 3.21 (0.07) 1.18 (0.01) 0.61 (0.01) MHMM-G 7.08 (0.07) 32.23 (0.06) 2.57 (0.17) 1.27 (0.03) 0.74 (0.02) MHMM-$\beta$ 1.28 (0.08) 4.37 (0.22) 2.39 (0.19) 9.09 (0.48) 0.67 (0.03) 0.76 (0.02) MHMM-$\beta^{*}$ 500 0.51 (0.03) 1.88 (0.08) 0.00 (0.00) 2.87 (0.13) 0.08 (0.00) HMM-G 6.55 (0.02) 32.47 (0.03) 1.14 (0.08) 0.89 (0.02) 0.89 (0.01) HMM-$\beta$ 2.81 (0.07) 9.91 (0.17) 2.35 (0.12) 0.92 (0.03) 0.72 (0.01) MHMM-G 7.09 (0.05) 32.03 (0.05) 1.71 (0.17) 1.10 (0.03) 0.82 (0.02) MHMM-$\beta$ 1.04 (0.07) 3.33 (0.19) 2.16 (0.19) 7.60 (0.41) 0.61 (0.04) 0.78 (0.02) MHMM-$\beta^{*}$ 1000 0.35 (0.01) 1.41 (0.05) 0.00 (0.00) 1.98 (0.08) 0.06 (0.00) HMM-G 6.83 (0.02) 31.93 (0.02) 0.82 (0.04) 0.85 (0.01) 0.93 (0.00) HMM-$\beta$ 2.03 (0.04) 9.08 (0.13) 0.37 (0.06) 0.26 (0.01) 0.96 (0.01) MHMM-G 6.98 (0.05) 32.04 (0.04) 1.57 (0.18) 1.09 (0.03) 0.85 (0.02) MHMM-$\beta$ 0.77 (0.05) 2.56 (0.15) 0.80 (0.15) 7.19 (0.34) 0.31 (0.03) 0.92 (0.02) MHMM-$\beta^{*}$ 2000 0.24 (0.01) 0.93 (0.04) 0.00 (0.00) 1.40 (0.06) 0.04 (0.00) HMM-G 6.97 (0.01) 31.67 (0.01) 0.48 (0.03) 0.87 (0.01) 0.97 (0.00) HMM-$\beta$ 1.65 (0.02) 8.04 (0.10) 0.06 (0.01) 0.20 (0.00) 1.00 (0.00) MHMM-G 6.87 (0.05) 32.04 (0.04) 0.79 (0.16) 1.02 (0.02) 0.92 (0.02) MHMM-$\beta$ 0.77 (0.07) 2.50 (0.21) 0.30 (0.09) 6.83 (0.38) 0.23 (0.03) 0.97 (0.01)
5 Case Study: Integrative Analysis of Quantum Dots
@@ 5.1 Cluster Patterns
We applied the proposed MHMM-$\beta$ approach with varying number of clusters. The cluster patterns are displayed in Figure 4, in which the colors or the gray levels represent different clusters and the stripes show how the dots are put into more and more clusters from the left to the right.
@@
Starting from the 2-cluster model, it can be seen that in the 3-cluster model, the third cluster is formed by QDs from both clusters. From there, as the number of clusters increases, the new clusters are mainly formed from further splitting the third cluster, and the first two clusters remain relatively stable throughout. This clear pattern suggests that the 3-cluster model is the most stable and informative. We have also computed AIC, BIC and ICL values for these models and unfortunately they have large discrepancy: AIC and BIC would select a model with more than 3 clusters while ICL suggests 2 clusters. This in fact is consistent with our simulation studies and comparative studies in the literature, which have shown that when the clusters are too overlapped, ICL tends to favor less number of clusters than BIC. From these results and after consulting with our chemistry collaborators, we decided to mainly focus on the results from the three-cluster model; as to be shown below, the results indeed are highly interpretable based on principles of chemistry and physics. We also provide the results from other models in Section D of the Supplementary Material.
5.2 Results from 3-Cluster MHMM-$\beta$ Model
In Figure 5, the right panel shows the estimated 0/1 inflated Beta distributions of the three intensity states from the MHMM-$\beta$ model. With the estimated state series, we obtain the empirical distributions of the three states directly from the standardized intensity data, which are shown in the left panel. As expected, the estimated and the empirical distributions are quite similar to each other, suggesting that here are no apparent outliers, and the MHMM-$\beta$ model fits the data well and successfully captures the collective behaviors of intensity fluctuations of all QDs.